OpenStudy (anonymous):

derivative application.... Water is flowing into a spherical tank with 6 ft radius at constant rate of 30pi cu ft per hr. When the water is h feet deep, the volume of water is given by V=((pih^2)/3)*(18-h) What is the rate at which the depth of the water in tank is increasing at the moment when the water is 2 ft deep?

OpenStudy (dumbcow):

\[\frac{dV}{dt} = \frac{dV}{dh}*\frac{dh}{dt}\] \[\rightarrow \frac{dh}{dt} = \frac{30}{dV/dh}\] take derivative and then plug in h=2

OpenStudy (anonymous):

Can you please explain how you would take the derivative?

OpenStudy (dumbcow):

\[V = \frac{\pi}{3}h^2 (18-h) = 6 \pi h^2 - \frac{\pi}{3} h^3\] you take derivative of polynomial \[\frac{d}{dx} x^n = n x ^{n-1}\]

OpenStudy (anonymous):

Thanks so much!

OpenStudy (dumbcow):

yw :)